OpenStudy (anonymous):

What is the 5th partial sum of Infinity sigma n=1 -2+5n

OpenStudy (agreene):

\[\sum_{n=1}^{m}(5n-2)=\frac12(5m^2+m)\] \[\therefore \frac12(5(5^2)+5)=65\] Alternatively... you can do it the long way: \[(5(1)-2)+(5(2)-2)+(5(3)-2)+(5(4)-2)+(5(5)-2)=65\]